The sum of the ages of Leah and Angel is 37. Three years ago, the difference of their ages was 17. How old is Leah?
Answer by Edwin McCravy(20055)

Let the older female's age by x and the youger female's age by y.

>>...The sum of the ages of Leah and Angel is 37...<<
x + y = 37

>>... Three years ago, the difference of their ages was 17...<<
Three years ago, the older female was x-3 and the younger female 
was y-3

So (x-3) - (y-3) = 17

So we have this system of equations:

        x + y = 37
(x-3) - (y-3) = 17

Simplify the second and solve the system and you'll get x = 27, y = 10.

>>...How old is Leah?...<<
The problem does not state which female is named Leah.  If Leah is
the older, then she is 27 and Angel is 10.  If Angel is the older then
she is 27 and Leah is 10.

Checking:
The sum of their ages is 27+10 or 37.  That checks.
Three years ago the older was 24 and the younger was 7, and 24-7 = 17.
That checks.  So the answers are right.  We just don't know which one is
named Leah.

Edwin